OpenStudy (anonymous):

Write the following equation in the general form Ax + By + C = 0. 1/2y - 1/3 x - 1 = 0 -2x + 3y - 6 = 0 2x - 3y + 6 = 0 2x - 3y - 6 = 0

OpenStudy (unklerhaukus):

none of the above

OpenStudy (anonymous):

it has to be one of them?

OpenStudy (unklerhaukus):

y - x - 1 = 0 -y + x + 1 = 0 x - y + 1 = 0 2x -2y + 2 = 0

OpenStudy (anonymous):

Write the following equation in the general form Ax + By + C = 0. 1/2y - 1/3 x - 1 = 0 -2x + 3y - 6 = 0 2x - 3y + 6 = 0 2x - 3y - 6 = 0

OpenStudy (unklerhaukus):

ah that's a bit different

OpenStudy (unklerhaukus):

1/2y - 1/3 x - 1 = 0 to get rid of the fractions , multiply all terms by 6

OpenStudy (anonymous):

im so confused

OpenStudy (unklerhaukus):

\[\tfrac12y - \tfrac13 x - 1 = 0\] multiply both sides of the equation by 6\[6\times(\tfrac12y - \tfrac13 x - 1) = 6\times0\] distributing \[(6\times\tfrac12y) + (6\times-\tfrac13 x) +(6\times- 1) = 6\times0\] simplify...